OpenStudy (anonymous):

A rock is suspended from a scale reads 20.0N. A beaker of water is raised up so the rock is totally submerged in the water. The scale now reads 12.5N. What is the density of the rock?

OpenStudy (anonymous):

Hmmm I would think that we would need to know the volume of the rock? What does someone else think?

OpenStudy (anonymous):

DISCLAIMER! I have pilfered this answer from a forum on the internet somewhere. Thank you neal448, whoever you are! "The mass of the rock is 20.0 N divided by the gravitational constant, 9.81 m/s2. Mass = 2.039 kg. While submerged in water, the rock has displaced (20.0 N - 12.5 N =) 7.5 N worth of water, which, when converted to mass equals 0.7645 kg. The density of water is 1000 kg/m3, so the volume of water displaced is (0.7645 kg) / (1000 kg/m3) = 7.645 x 10^-4 m3. The rock's density = mass(rock) / volume(rock) = (2.039 kg) / (7.645 x 10^-4 m3) = 2667 kg/m3."

OpenStudy (anonymous):

yep i fully agree with you walking stick